# Mitigating Cache Noise in Test-Time Adaptation for Large Vision-Language Models

### v1

@@ 1 Introduction
To address these limitations, we propose the “Cache, Residual, Gaussian” (CRG) zero-shot test-time adaptation method, which enhances cache reliability via a comprehensive mechanism. CRG maintains two caches for positive and negative image features, plus a cache for text prototypes. Image prototypes are derived by averaging features within each class, and learnable residual vectors are used to align text and image prototypes by minimizing prediction entropy and maximizing inter-prototype distance. However, noisy samples may persist in the cache. To mitigate their impact, CRG employs Gaussian Discriminant Analysis (GDA) and Bayes’ theorem to model class distributions and compute posterior probabilities. This distribution-based decision-making reduces noise interference, while negative class prototypes further curb overconfidence caused by corrupted features. Through the synergy of GDA and negative prototypes, CRG achieves robust, accurate performance even under noisy conditions.
@@ 2 Related Works
2.1 Vision-Language Models
Recent years have witnessed rapid progress in foundational models, particularly vision-language models, due to their impressive generalization capabilities. Specifically, CLIP aligns vision and text embeddings through contrastive learning on large-scale, diverse image-text datasets. However, due to the distribution shift between training and downstream data, many researchers have explored fine-tuning CLIP with a small amount of labeled data to address distribution shift and improve adaptation. Existing CLIP fine-tuning approaches are mainly categorized into two types[16]: prompt-based methods (e.g., CoOp[31] and Maple[14]) and adapter-based methods (e.g., Tip-Adapter[30] and CLIP-Adapter[8]). Nonetheless, these few-shot fine-tuning methods are limited to the distribution of a small number of labeled samples, restricting their generalization to broader distributions. In this study, we focus on test-time adaptation under zero-shot settings, as acquiring labeled samples in real-world deployments is often impractical.
@@ 2.2 Test-Time Adaptation for Vision-Language Models
Recently, test-time adaptation of large-scale vision-language models has emerged as a research hotspot. TPT[24] (Test-Time Prompt Tuning) integrates consistency regularization with prompt tuning to enforce consistency between augmented views of test samples. DiffTPT[7] improves upon TPT with stronger augmentations but incurs high computational costs. PromptAlign[10] aligns visual and textual features using statistical characteristics of a proxy dataset. TDA[13] introduces a lightweight cache to store low-entropy historical test features to guide classification. However, these methods often require gradient backpropagation or remain vulnerable to noise in cached features.To address these challenges, we design CRG. CRG aligns prototypes by learning residuals without requiring backpropagation through the encoder. By integrating Gaussian Discriminant Analysis[2] (GDA) and negative learning, it achieves greater robustness under noisy label conditions.
@@ 3 Method
In this section, we first provide a brief overview of the foundational concepts of CLIP[22]. We then present our method by following the sequence of Cache, Residual, and Gaussian, outlining each component in detail to highlight how they contribute to the overall framework.
@@ 3.1 Preliminaries
Contrastive Language-Image Pre-training (CLIP). CLIP[22] utilizes two pre-trained parallel encoders: a visual encoder $E_{V}(\cdot)$ and a textual encoder $E_{T}(\cdot)$, which embed images and text descriptions into a shared embedding space $\mathbb{R}^{d}$. For a $C$-class classification task, CLIP performs zero-shot predictions by computing the similarities between the extracted image feature and the $C$ candidate text features, formulated as:
$P_{\text{CLIP}}(y=c\mid X_{\text{test}})=\frac{\exp\left(\text{sim}\left(E_{T}(T_{c}),E_{V}(X_{\text{test}})\right)/\tau\right)}{\sum_{j=1}^{C}\exp\left(\text{sim}\left(E_{T}(T_{j}),E_{V}(X_{\text{test}})\right)/\tau\right)},$ (1)
where $X_{\text{test}}\in D_{\text{test}}$ denotes the input test image, and $T_{c}$ represents the class-specific text description for class $y_{c}$. The pairwise similarity $\text{sim}(\cdot,\cdot)$ is calculated using cosine similarity, and $\tau$ is the temperature parameter in the softmax function.
@@
3.2 Cache
Inspired by the method of TDA[13], we adopt a priority queue strategy to store image features for each class. Specifically, for each test image $X_{\text{test}}$, after making a prediction, we assign it to the corresponding class queue based on its predicted pseudo-label $\hat{y}$. Meanwhile, we compute the entropy value $h_{\text{test}}$ of the test image and store the image feature together with its entropy as a pair $({E_{V}}(X_{\text{test}}),h_{\text{test}})$ in the queue. For each class $c$, the size of its priority queue is $\mathcal{M}$, and each queue is initialized with a text feature $t_{c}$, the reasoning for which will be detailed in the Gaussian section.
When the queue is full, we compare the entropy value $h_{\text{test}}$ of the new test image with the highest entropy value in the queue: if the new image has higher entropy, it will be discarded; if the entropy is lower, the image with the highest entropy in the queue will be replaced by the new image. If the queue is not yet full, the new image is inserted directly.
We also construct a text-side cache $T_{\text{cache}}$ to store class-specific text features derived from text descriptions. The cache has the shape $T_{\text{cache}}\in\mathbb{R}^{C\times d}$, where
$T_{\text{cache}}=[t_{1},t_{2},\dots,t_{C}],$
and each $t_{c}$ represents the text feature for class $c$.
To incorporate distribution information from historical samples, $T_{\text{cache}}$ is continuously updated during inference. Specifically, let $T_{\text{cache}}^{*}$ be the text cache optimized for each test sample in Equation 2. To retain historical information, we update the text-side over time like DPE[29]
$T_{\text{cache}}\leftarrow\frac{(k-1)T_{\text{cache}}+T_{\text{cache}}^{*}}{\|(k-1)T_{\text{cache}}+T_{\text{cache}}^{*}\|},\quad k\leftarrow k+1$
To ensure stable online updates, we set an entropy threshold $\tau_{t}$ to filter out low-confidence samples (for which $H(P(X_{\text{test}}))<\tau_{t}$) from updating the online prototypes, and maintain a counter $k$ for tracking confident samples.
3.3 Residual
We observe that residual learning holds significant potential [26] for improving model performance. More importantly, we introduce learnable residuals on both the text and vision sides, enabling feature space calibration for each test sample during inference. During testing, we learn three types of residuals: text cache residuals and vision cache residuals, with the latter further divided into positive and negative cache residuals.
Text Cache Residuals. We introduce learnable residual parameters $R_{T}$ to dynamically adjust the text cache during testing. Specifically, the text cache is updated as follows:
${T}_{\text{cache}}=\text{Normalize}({T}_{\text{cache}}+R_{T})$ (2)
Here, $R_{T}\in\mathbb{R}^{C\times d}$ is a learnable parameter initialized to zero.
Positive Cache Residuals. For each class $c$, we compute the average of the features stored in the visual cache to obtain a cache prototype representing each class, denoted as $V_{\text{cache}}^{+}=[v_{1}^{+},v_{2}^{+},\dots,v_{C}^{+}]\in\mathbb{R}^{C\times d}$. Subsequently, we introduce a set of visual residuals $R_{V}^{+}$ to further refine these prototypes. Specifically, the updated visual prototype is computed as:
$V_{\text{cache}}^{+}=\text{Normalize}(V_{\text{cache}}^{+}+R_{V}^{+})$ (3)
where $R_{V}^{+}\in\mathbb{R}^{C\times d}$ is the learnable residual parameter, initialized to zero.
Negative Cache Residuals. Unlike TDA[13], which treats high-entropy samples as negative samples, our approach leverages the principle that if an image is classified as a dog, it cannot simultaneously be a cat. Therefore, instead of constructing a separate negative sample cache, we mine the negative prototype directly from the existing cache.
To construct the negative cache prototypes $V_{\text{cache}}^{-}=[v_{1}^{-},v_{2}^{-},\dots,v_{n}^{-}]\in\mathbb{R}^{C\times d}$, for each class $c$, we compute the average of the prototypes from all other classes, effectively excluding class $c$. Formally, the negative prototype is defined as:
${v}_{c}^{-}=\frac{1}{C-1}\sum_{\begin{subarray}{c}j=1\\ j\neq c\end{subarray}}^{C}v_{j}^{+}$ (4)
To further refine these negative prototypes, we introduce a set of learnable residual parameters $R_{V}^{-}\in\mathbb{R}^{C\times d}$. The updated negative prototypes are computed as:
where $V_{\text{cache}}^{-}$ is the set of the negtive protptypes and $R_{V}^{-}\in\mathbb{R}^{C\times d}$ is the learnable residual parameter, initialized to zero.
@@ 3.3 Residual
The learning diagram of CRG can be illustrated by Fig.2. Given the complexity of Gaussian Discriminant Analysis and the risk of instability in high-dimensional spaces, we continue to use the similarity matching method to adjust class prototypes. Using the text cache, positive cache prototypes, and negative cache prototypes, the prediction for the input sample $X$ is as follows:
$P(y=c\mid X)=\frac{\exp\left((f_{v}^{\top}t_{c}+\mathcal{A}(f_{v}^{\top}v_{c}^{+})+\mathcal{B}(f_{v}^{\top}v_{c}^{-}))/\tau\right)}{\sum_{j=1}^{C}\exp\left((f_{v}^{\top}t_{j}+\mathcal{A}(f_{v}^{\top}v_{j}^{+})+\mathcal{B}(f_{v}^{\top}v_{j}^{-}))/\tau\right)}$ (6)
@@
Similar to the loss function used in TPT[24], we optimize these three residuals to promote consistent predictions across $N$ different augmented views of the given test image $X_{\text{test}}$ using the unsupervised entropy minimization objective.
$\mathcal{L}_{\text{TPT}}=H\left(\frac{1}{\rho N}\sum_{i=1}^{N}\mathbb{I}\left(H(P(y\mid\tilde{X}_{i}))<\theta\right)P(y\mid\tilde{X}_{i})\right),$ (7)
where $P(y\mid\tilde{X}_{i})$ is the predicted probability distribution for the augmented view $\tilde{X}_{i}$, $H(\cdot)$ is the entropy function, $\theta$ is an entropy threshold, $\rho$ represents the filtered ratio, and $\mathbb{I}(\cdot)$ is the indicator function ensuring that only augmented views with low entropy are used for training.
@@
$\mathcal{L}_{\text{positive}}^{\text{negative}}=\sum_{c=1}^{C}\frac{\langle v_{c}^{+},v_{c}^{-}\rangle}{\|v_{c}^{+}\|\|v_{c}^{-}\|}.$ (9)
Here, Equation 8 is used to separate the text prototype, while Equation 9 is used to separate the positive cache prototype and the negative cache prototype, where $t_{m}$ is the text features of the $m$-th class, $v_{c}^{+}$ denotes the positive cache prototype, and $v_{c}^{-}$ denotes the negative cache prototype.
@@
Through these approaches, we can achieve better alignment of the three different prototype features.
@@ 3.4 Gaussian
Theoretical Justification. Recent studies[17] have theoretically demonstrated that features follow a Gaussian distribution when the network is trained with the Softmax function. Since CLIP uses the Softmax function during training, this provides a theoretical justification for applying Gaussian Discriminant Analysis[2]. In previous calibration processes, learning a residual essentially corresponds to applying a simple linear transformation to the features. If the original features follow a Gaussian distribution, the calibrated features will also retain the Gaussian distribution property.
Gaussian Discriminant Analysis. We first introduce the assumption of GDA[2], where the features $x$ under class $i$ follow a Gaussian distribution. Specifically, the class-conditional distribution can be expressed as:
$p(x\mid y=i)\sim\mathcal{N}(\mu_{i},\Sigma),$
where $\mu_{i}$ is the mean vector for class $i$, and $\Sigma$ is the shared covariance matrix across all classes.
Based on this assumption, we use Bayes’ theorem to compute the posterior probability $p(y=i\mid x)$, given by:
$p(y=i\mid x)=\frac{p(x\mid y=i)p(y=i)}{\sum_{j=1}^{C}p(x\mid y=j)p(y=j)}.$
Substituting the Gaussian form of $p(x\mid y=i)\sim\mathcal{N}(\mu_{i},\Sigma),$ into the above equation, the posterior probability can be further written as follows. The detailed derivation process is provided in the appendix. :
$p(y=i\mid x)=\frac{\exp\left(\mu_{i}^{T}\Sigma^{-1}x-\frac{1}{2}\mu_{i}^{T}\Sigma^{-1}\mu_{i}+\log p_{i}\right)}{\sum_{j=1}^{C}\exp\left(\mu_{j}^{T}\Sigma^{-1}x-\frac{1}{2}\mu_{j}^{T}\Sigma^{-1}\mu_{j}+\log p_{j}\right)},$
where $p_{i}=p(y=i)=\frac{1}{C}$ for $i=1,2,\dots,C$, assuming a uniform prior distribution across all classes.
Through normalization of the posterior probability, the classifier can be equivalently represented as a linear classifier, where the weight $w_{i}$ and bias $b_{i}$ are defined as:
$w_{i}=\Sigma^{-1}\mu_{i},\quad b_{i}=\log p_{i}-\frac{1}{2}\mu_{i}^{T}\Sigma^{-1}\mu_{i}.$
@@
${h}_{i}(x)=w_{i}^{T}x+b_{i}.$ (11)
In the actual inference process, we integrate the learned residuals into the cached features of each category to obtain well-aligned positive cache features. We then compute the mean for each class and the overall covariance to construct a Gaussian discriminant classifier, which replaces the original positive sample inference method. Additionally, to prevent some category queues from being empty at initialization, we insert a pair consisting of a textual vector and a high-entropy value into each category queue.
Finally, for the final inference, we replace $\mathcal{A}(f_{v}^{\top}v_{c}^{+})$ with $\lambda_{1}h_{c}(f_{v})$. Following the format of Equations (6) and (11), this ultimately leads to Equation (12).
$P(y=c\mid X)=\frac{\exp\left((f_{v}^{\top}t_{c}+\lambda_{1}{h}_{c}(f_{v})+\mathcal{B}(f_{v}^{\top}v_{c}^{-}))/\tau\right)}{\sum_{j=1}^{C}\exp\left((f_{v}^{\top}t_{j}+\lambda_{1}{h}_{i}(f_{v})+\mathcal{B}(f_{v}^{\top}v_{j}^{-}))/\tau\right)}$ (12)

### v2

@@ 1 Introduction
We introduce the Cache, Residual, Gaussian” (CRG) zero-shot test-time adaptation method to overcome these limitations, which features a comprehensive mechanism to enhance cache reliability. It includes two caches that separately store the positive and negative feature sets of image samples, along with an additional cache dedicated to storing text prototypes[26]. The prototypes for the image features are calculated by averaging the features within the same class[35]. To achieve thorough alignment between text and image prototypes, CRG employs learnable residual vectors [35; 26], which are optimized by minimizing prediction entropy[24] and maximizing inter-prototype distance.
Despite achieving modality alignment, the cache may still harbor noisy features. To mitigate their effect, we model class distributions using the Gaussian Discriminant Analysis (GDA) framework [2; 28] and utilize Bayes’ theorem to calculate posterior probabilities. This decision-making approach, grounded in distribution, effectively minimizes the impact of noisy samples. Additionally, to reduce the overconfidence in predictions induced by noisy features, we computed negative class prototypes, which serve as counter-references to further mitigate noise interference. Owing to the synergistic interaction of GDA and negative prototypes, CRG sustains heightened robustness and accuracy even in the presence of noise.
@@ 2.1 Test-Time Adaptation for Vision-Language Models
Due to distribution shifts between training and downstream data, many researchers have explored fine-tuning CLIP with a small amount of labeled data to mitigate these shifts and enhance adaptability. Existing CLIP fine-tuning methods can be divided into two main categories[16]: prompt-based methods (e.g., CoOp[38] and MaPLe[14]) and adapter-based methods (e.g., Tip-Adapter[36] and CLIP-Adapter[8]). Recently, test-time adaptation for large-scale vision-language models has attracted significant interest. TPT[24] (Test-time Prompt Tuning) combines consistency regularization with prompt tuning to reinforce consistency among augmented views of test samples. DiffTPT[7] builds on TPT, offering stronger augmentation capabilities at a higher computational cost. PromptAlign[10] employs the statistical features of a proxy dataset to align visual and textual representations. DPE[35] introduces visual-textual prototype evolution to optimize modality alignment, while DMN[37] utilizes a dynamic cache to store test-time prototype memories. TDA[13] further enhances performance by incorporating a negative visual cache in addition to the positive visual cache.
2.2 Negative Prototype Learning
Prototype learning[30] methods represent classes with prototypes and perform classification based on similarities to these prototypes. Many prototype learning approaches focus solely on visual modality tasks; for instance, DPNP[33] introduces a discriminative model built upon deep positive and negative prototypes. Decoupled Prototype[27] Learning applies positive and negative prototypes to online test-time adaptation, reducing performance degradation caused by noisy pseudo-labels. In visual-language models, SimNL[34] enhances model adaptability in few-shot scenarios by constructing negative features in both the textual and visual modalities. In contrast, this paper introduces, for the first time, the concept of negative visual prototype learning for visual-language model test-time adaptation tasks. It aims to better align visual and textual prototypes in zero-shot scenarios by exploiting negative information at test time.
@@ 3 Method
Our work builds upon the CLIP visual-language model by incorporating the prototype residual vector learning technique from DPE[35]. We innovatively propose a reliable Test-Time Adaptation (TTA) framework that includes caches, three sets of prototype residual learning, and decision-making based on Gaussian discriminant analysis.
@@ 3.1 Preliminaries
CLIP[22] (Contrastive Language-Image Pretraining) maps images and texts into a shared embedding space using an image encoder $f$ and a text encoder $g$. In the zero-shot classification scenario, for an image $I$ and candidate classes $\{c_{1},c_{2},\ldots,c_{K}\}$ with associated text prompts $\{T_{1},T_{2},\ldots,T_{K}\}$, the probability of the image belonging to class $c_{k}$ is computed as follows:
$p(c_{k}\mid I)=\frac{\exp\left(\tau\cdot\operatorname{sim}(f(I),g(T_{k}))\right)}{\sum_{j=1}^{K}\exp\left(\tau\cdot\operatorname{sim}(f(I),g(T_{j}))\right)}$
Here, the cosine similarity is defined as $\operatorname{sim}(x,y)=\frac{x\cdot y}{\|x\|\|y\|}$, where: $\tau$ is a temperature scaling parameter, $K$ is the number of classes.
3.2 Caches
Inspired by the method of TDA[13], we adopt a priority queue strategy to store image features for each class. Specifically, for each test image $I$, after making a prediction $p(c_{k}\mid I)$, we assign it to the corresponding class queue based on its predicted pseudo-label $\hat{y}$. Meanwhile, we compute the entropy value $h_{I}$ of the test image and store the image feature together with its entropy as a pair $(f(I),h_{I})$ in the queue. For each class $c_{k}$, the size of its priority queue is $\mathcal{M}$, and each queue is initialized with a text feature $g(T_{k})$, the reasoning for which will be detailed in the Gaussian section.
When the queue is full, we compare the entropy value $h_{I}$ of the new test image with the highest entropy value in the queue: if the new image has higher entropy, it will be discarded; if the entropy is lower, the image with the highest entropy in the queue will be replaced by the new image. The new image is inserted directly if the queue is not yet full.
We also construct a text-side cache $T_{\text{cache}}$ to store class-specific text features derived from text descriptions. The cache has the shape $T_{\text{cache}}\in\mathbb{R}^{K\times d}$, where
$T_{\text{cache}}=[t_{1},t_{2},\dots,t_{K}],$
and each $t_{k}=g(T_{k})$ represents the text feature for class $c_{k}$.
We update the text cache during inference to capture historical distribution information. We adopt DPE[35]’s momentum update, using only high-confidence samples (entropy above threshold) to maintain stability.
3.3 Residual
Residual learning shows significant potential for enhancing model performance[26; 35; 32]. More notably, we introduce learnable prototype residuals across the language, positive visual, and negative visual domains, enabling feature space calibration for each test sample during inference.
Text Cache Residuals. We use learnable residual parameters $R_{T}$ to dynamically adjust the text cache during testing. Specifically, the text cache is updated as follows:
${T}_{\text{cache}}=\text{Normalize}({T}_{\text{cache}}+R_{T})$ (1)
Here, $R_{T}\in\mathbb{R}^{K\times d}$ is a learnable parameter initialized to zero.
Positive Cache Residuals. For each class $c_{k}$, we compute the average of the features stored in the visual cache to obtain a cache prototype representing each class like DPE[35], denoted as $V_{\text{cache}}^{+}=[v_{1}^{+},v_{2}^{+},\dots,v_{K}^{+}]\in\mathbb{R}^{K\times d}$. Subsequently, we use a set of visual residuals $R_{V}^{+}$ to further refine these prototypes. Specifically, the updated visual prototype is computed as:
$V_{\text{cache}}^{+}=\text{Normalize}(V_{\text{cache}}^{+}+R_{V}^{+})$ (2)
where $R_{V}^{+}\in\mathbb{R}^{K\times d}$ is the learnable residual parameter, initialized to zero.
Negative Cache Residuals. Unlike TDA[13], which treats high-entropy samples as negative samples, our approach is based on the intuitive principle that if an image is classified as a dog, it cannot simultaneously be a cat—a concept inspired by SimNL[34]. However, unlike both SimNL[34] and TDA[13], we do not build a separate negative sample cache; instead, we directly extract negative prototypes from the existing positive visual prototypes, our method can be conceptually regarded as having a “virtual” negative cache.
To construct the negative cache prototypes $V_{\text{cache}}^{-}=[v_{1}^{-},v_{2}^{-},\dots,v_{n}^{-}]\in\mathbb{R}^{K\times d}$, for each class $c_{k}$, we compute the average of the prototypes from all other classes, effectively excluding class $c_{k}$. Formally, the negative prototype is defined as:
${v}_{k}^{-}=\frac{1}{K-1}\sum_{\begin{subarray}{c}j=1\\ j\neq k\end{subarray}}^{K}v_{j}^{+}$ (3)
To further refine these negative prototypes, we introduce a set of learnable residual parameters $R_{V}^{-}\in\mathbb{R}^{K\times d}$. The updated negative prototypes are computed as:
where $V_{\text{cache}}^{-}$ is the set of the negtive protptypes and $R_{V}^{-}\in\mathbb{R}^{K\times d}$ is the learnable residual parameter, initialized to zero.
@@ 3.3 Residual
The learning diagram of CRG can be illustrated by Fig.2. Given the complexity of Gaussian Discriminant Analysis and the risk of instability in high-dimensional spaces, we continue to use the similarity matching method to adjust class prototypes. Using the text cache, positive cache prototypes, and negative cache prototypes, the prediction for the input sample $I$ is as follows:
$P(c_{k}\mid I)=\frac{\exp\left((f_{I}^{\top}t_{k}+\mathcal{A}(f_{I}^{\top}v_{k}^{+})+\mathcal{B}(f_{I}^{\top}v_{k}^{-}))/\tau\right)}{\sum_{j=1}^{K}\exp\left((f_{I}^{\top}t_{j}+\mathcal{A}(f_{I}^{\top}v_{j}^{+})+\mathcal{B}(f_{I}^{\top}v_{j}^{-}))/\tau\right)}$ (5)
@@
Similar to the loss function used in TPT[24], we optimize these three residuals to promote consistent predictions across $N$ different augmented views of the given test image $I$ using the unsupervised entropy minimization objective.
$\mathcal{L}_{\text{TPT}}=H\left(\frac{1}{\rho N}\sum_{i=1}^{N}\mathbb{I}\left(H(p(y\mid\tilde{I}_{i}))\leq\theta\right)P(y\mid\tilde{I}_{i})\right),$ (6)
where $p(y\mid\tilde{I}_{i})$ is the predicted probability distribution for the augmented view $\tilde{I}_{i}$, $H(\cdot)$ is the entropy function, $\theta$ is an entropy threshold, $\rho$ represents the filtered ratio, and $\mathbb{I}(\cdot)$ is the indicator function ensuring that only augmented views with low entropy are used for training.
@@
$\mathcal{L}_{\text{positive}}^{\text{negative}}=\sum_{c=1}^{K}\frac{\langle v_{c}^{+},v_{c}^{-}\rangle}{\|v_{c}^{+}\|\|v_{c}^{-}\|}.$ (8)
Here, Equation 7 is used to separate the text prototype. In contrast, Equation 8 is used to separate the positive cache prototype and the negative cache prototype, where $t_{m}$ is the text features of the $m$-th class, $v_{c}^{+}$ denotes the positive cache prototype, and $v_{c}^{-}$ denotes the negative cache prototype.
@@
where $\xi_{1}$ and $\xi_{2}$ are two weight hyperparamenters. We can achieve better alignments of the three different prototype features through these approaches.
@@ 3.4 Gaussian
Theoretical Justification. Recent studies[17] have theoretically demonstrated that features follow a Gaussian distribution when the network is trained with the Softmax function. Since CLIP uses the Softmax function during training, this provides a theoretical justification for applying Gaussian Discriminant Analysis[2]. In previous calibration processes, learning a residual essentially corresponds to applying a simple linear transformation to the features. If the original features follow a Gaussian distribution, the calibrated features will also retain the Gaussian distribution property. Furthermore, [28] has successfully applied GDA in few-shot, base-to-novel, and unsupervised scenarios to adapt visual-language models.
Gaussian Discriminant Analysis. We first introduce the assumption of GDA, where the feature $f_{I}$ under class $k$ follow a Gaussian distribution. Specifically, the class-conditional distribution can be expressed as:
$p(I\mid y=k)\sim\mathcal{N}(\mu_{k},\Sigma),$
where $\mu_{k}$ is the mean vector for class $c_{k}$, and $\Sigma$ is the shared covariance matrix across all classes.
Based on this assumption, we use Bayes’ theorem to compute the posterior probability $p(y=k\mid I)$, given by:
$p(y=k\mid I)=\frac{p(f_{I}\mid y=k)p(y=k)}{\sum_{j=1}^{K}p(f_{I}\mid y=j)p(y=j)}.$
Substituting the Gaussian form of $p(I\mid y=k)\sim\mathcal{N}(\mu_{k},\Sigma),$ into the above equation, the posterior probability can be further written as follows. The detailed derivation process is provided in the appendix. :
$p(y=k\mid I)=\frac{\exp\left(\mu_{k}^{T}\Sigma^{-1}f_{I}-\frac{1}{2}\mu_{k}^{T}\Sigma^{-1}\mu_{k}+\log p_{k}\right)}{\sum_{j=1}^{K}\exp\left(\mu_{j}^{T}\Sigma^{-1}f_{I}-\frac{1}{2}\mu_{j}^{T}\Sigma^{-1}\mu_{j}+\log p_{j}\right)},$
where $p_{k}=p(y=k)=\frac{1}{K}$ for $k=1,2,\dots,K$, assuming a uniform prior distribution across all classes.
Through normalization of the posterior probability, the classifier can be equivalently represented as a linear classifier, where the weight $w_{k}$ and bias $b_{k}$ are defined as:
$w_{k}=\Sigma^{-1}\mu_{k},\quad b_{k}=\log p_{k}-\frac{1}{2}\mu_{k}^{T}\Sigma^{-1}\mu_{k}.$
@@
${h}_{k}(f_{I})=w_{k}^{T}f_{I}+b_{k}.$ (10)
In the actual inference process, we use the well-aligned positive cache features to compute the mean for each class and the overall covariance, thereby constructing a Gaussian discriminant classifier. Additionally, to prevent some category queues from being empty at initialization, we insert a pair consisting of a textual vector and a high-entropy value into the queue of each category.
Finally, when providing the final inference results, we replace $\mathcal{A}(f_{I}^{\top}v_{k}^{+})$ with $\lambda_{1}{h}_{k}(f_{I})$, following the format of Equation 5 and 10, we get Equation 11.
$p(y=k\mid I)=\frac{\exp\left((f_{I}^{\top}t_{k}+\lambda_{1}{h}_{k}(f_{I})+\mathcal{B}(f_{I}^{\top}v_{k}^{-}))/\tau\right)}{\sum_{j=1}^{K}\exp\left((f_{I}^{\top}t_{j}+\lambda_{1}{h}_{i}(f_{I})+\mathcal{B}(f_{I}^{\top}v_{j}^{-}))/\tau\right)}$ (11)